) If f(x) = 5x^7 - 3, then f^-1(x) =
a. [(x + 7)/5]^1/3
b. [(x + 3)/5]^1/7
c. [(x + 5)/7]^1/3
d. [(x + 7)/3]^1/5
e. [(x + 3)/7]^1/5
f. [(x + 5)/3]^1/7
g. none of these

Answer by Nate(3500)

f(x) = 5x^7 - 3
y = 5x^7 - 3
x = 5y^7 - 3
x + 3 = 5y^7
(x + 3)/5 = y^7
([x + 3]/5)^(1/7) = f^-1(x)
answer: b.